Please solve this question.

Americans spend more than they make. A typical household spends $10,000 dollars more than it makes. They do this for 8 years. If this debt will be financed at 15% per year, what annual repayment will be required to repay the debt over a 10- year period?
Answer #1

Spending in excess of income is  $10,000 per year and this is done for 8 years so that the debt accumulated to

$80000 for 8 years.

This debt is now financed at 15% per year and 10 annual payments are required.

Find the annual payment using 80000(A/P, 15%, 10) = PMT(15%,10,-80000,0)

which results in a payment of 80000*0.19925 = $15,940

The formula for (A/P, 15%, 10) = (0.15*(1.15^10))/(1.15^10-1) = 0.19925.